OpenStudy (liizzyliizz):

domain and range of y=2^x ehh not sure about this one lol..

jimthompson5910 (jim_thompson5910):

You can plug in any number for x, so the domain is \[(-\infty,\infty)\]

OpenStudy (dumbcow):

b^x is always greater than zero for any positive base

jimthompson5910 (jim_thompson5910):

The outputs here are all positive, so the range is y > 0, which means that the range in interval notation is \[(0,\infty)\]